OpenStudy (anonymous):

I need help: Simplify 1/4(8m - 4n) +1/3 (6m + 3n).

OpenStudy (anonymous):

factor both denominators

OpenStudy (anonymous):

all it asks is to simplify, i don't know how to do it

OpenStudy (anonymous):

Multiply 4(8m-4n) and 3(6m+3n)

OpenStudy (paxpolaris):

i just wanted to make sure it look like the way i wrote it earlier .... and NOT: \[\frac 14 (8m-4n) + \frac 13(6m+3n)\]

OpenStudy (paxpolaris):

@jewjewbri ?

OpenStudy (paxpolaris):

@jewjewbri could you clarify, does your expression look like the first one or the second one?

OpenStudy (anonymous):

the second one

OpenStudy (paxpolaris):

then it is much easier: what is the common factor of (8m-4n)

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

a/c+b/d=ad+cb/cd

OpenStudy (paxpolaris):

\[\large 8m+4n= 2(4m+2n) = \dots\] keep going ... there is still some thing in common left

OpenStudy (anonymous):

If the question look like the first one.

OpenStudy (anonymous):

8m+4n=8m+2n

OpenStudy (paxpolaris):

we are simplifying:\[\Large \frac 14 (8m-4n) + \frac 13(6m+3n)\] all right @jewjewbri let's try this another way... Are you familiar with the distributive property?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Hey for the fist expression you can factor ou as (8m-4n)= 4(2m-n). It will be more simple

OpenStudy (paxpolaris):

\[\large a(b+c) = ab+ ac\]

OpenStudy (anonymous):

And the second, using the distribuitive property got (6m+3n)=3(2m+n)

OpenStudy (anonymous):

You got (6m+3n)=3(2m+n)

OpenStudy (anonymous):

Make some algebric now jewjewbri.

OpenStudy (anonymous):

6m+3n=6m+3n

OpenStudy (paxpolaris):

try it another way: \[\frac 14 \left( 8m-4n \right)=\frac14\cdot8m-\frac14\cdot 4n=...\] instead of doing one-fourth of the whole thing do 1 fourth of the parts...

OpenStudy (paxpolaris):

what is : \[\frac 14 \cdot 8m = {8m \over 4}=???\]

OpenStudy (anonymous):

Great tip.

OpenStudy (anonymous):

Just make the division of 8 by 4. Jewjewbri.

OpenStudy (paxpolaris):

@jewjewbri ?

OpenStudy (paxpolaris):

what is 8 divided by 4? @jewjewbri

OpenStudy (anonymous):

2

OpenStudy (paxpolaris):

so, 1-fourth of 8m is \(2m\)...

OpenStudy (paxpolaris):

same way 1-fourth of 4n is \(1n\), or just \(n\)

OpenStudy (paxpolaris):

so, \[\large \frac 14 \left( 8m-4n \right)=\color{green}{2m-n}\]

OpenStudy (paxpolaris):

you can do the second part the same way yourself: \[\large \frac 13 \left( 6n+3n \right)=\]

OpenStudy (paxpolaris):

2m+n